OpenStudy (anonymous):

What is the derivative of f(x) = [1+(x)^(1/2)][x^3]

OpenStudy (anonymous):

multiply in the x^3 then use power rule f(x) = x^3 + x^(7/2) f'(x) = 3x^2 + (7/2)x^(5/2)

OpenStudy (anonymous):

@jayz657 Thank you very much! Just curious, are you allowed to use the product rule if there is a sum/difference as one (or both) of the quantities, as in this problem?

OpenStudy (anonymous):

yea you can apply product rule for it, i just multiplied it in to make it easier